# ASURA-FDPS-ML: Star-by-star Galaxy Simulations Accelerated by Surrogate Modeling for Supernova Feedback

### v1

@@ Abstract
We introduce new high-resolution galaxy simulations accelerated by a surrogate model that reduces the computation cost by approximately 75 percent. Massive stars with a Zero Age Main Sequence mass of about 8 $\mathrm{M_{\odot}}$ and above explode as core-collapse supernovae (CCSNe), which play a critical role in galaxy formation. The energy released by CCSNe is essential for regulating star formation and driving feedback processes in the interstellar medium (ISM). However, the short integration timesteps required for SNe feedback present significant bottlenecks in star-by-star galaxy simulations, which aim to capture the individual stellar dynamics and the inhomogeneous shell expansion of SNe within the turbulent ISM. Our new framework combines direct numerical simulations and surrogate modeling, including machine learning and Gibbs sampling. The star formation history and the time evolution of outflow rates in the galaxy match those obtained from resolved direct numerical simulations. Our new approach achieves high-resolution fidelity while reducing computational costs, effectively bridging the physical scale gap and enabling multi-scale simulations.
@@
thanks: JSPS Research Fellowthanks: JSPS Research Fellow
@@ 1 Introduction
Supernovae (SNe) are powerful events that release an immense amount of energy into the ambient ISM ($\sim 10^{51}$ erg). This energy release results in immediate heating, and delayed momentum input into the ambient ISM drives turbulence and aids star formation regulation in the turbulent ISM (e.g., Somerville & Davé, 2015; Naab & Ostriker, 2017, for reviews). Thus, SNe significantly impact galaxy evolution, incorporating processes like gravity, hydrodynamics, radiation, star formation, and chemical reactions. Given the complicated interplay among these processes, numerical methods have been commonly employed to study galaxy formation and evolution.
Recently, numerical simulations have reached the resolution requirements to resolve three of the four phases of SN-remnant evolution from first principles with a typical mass resolution of around $\sim 1~{}M_{\odot}$ (e.g., Andersson et al., 2023, 2024; Deng et al., 2024; Emerick et al., 2018, 2019; Fujii et al., 2021, 2021, 2022b, 2022a, 2024; Gutcke et al., 2021, 2022; Hirai et al., 2021; Hu et al., 2016, 2017; Hu, 2019; Lahén et al., 2020, 2023; Smith et al., 2021; Steinwandel et al., 2020; Steinwandel & Goldberg, 2023; Steinwandel et al., 2023; Lahén et al., 2024; Steinwandel et al., 2024a, b). The three phases that can be resolved in these first principle simulations are the energy-conserving Sedov-Taylor phase, the pressure-driven snowplow phase, and the momentum-conserving snowplow phase (e.g., Sedov, 1959; Taylor, 1950; Blondin et al., 1998; von Neumann, 1942). These simulations resolve every phase of SN-remnant evolution apart from the phase of free expansion that is terminated when the swept-up mass becomes equal to the ejecta mass of the remnant of a few solar masses (e.g., Reynolds et al., 2008). This stage is when the reverse shock decouples from the forward shock, propagating backward into the vacuum generated by the forward shock. The reverse shock heats the gas inside the cavity generated by the forward shock to temperatures of $\sim 10^{7}$ K. This is why the injection of the SN-energy has been generally handled as a single thermal injection event so far, which can effectively be seen as a “subgrid model.”
To resolve this phase more accurately, one would need to have at least 10 to 100 times higher mass resolution than the ejecta mass, which requires switching from a thermal to a kinetic injection scheme. The resolved reverse shock would then ultimately heat the cavity generated by the forward shock and generate the SN-bubble that resolves the remnant into the well-studied energy conserving Sedov-Taylor-von-Neumann-solution (e.g., Kim & Ostriker, 2015; Steinwandel et al., 2020). Such a kinetic injection scheme has been applied in many successful simulation endeavors of the ISM (e.g., Hu et al., 2016), but this scheme does require not only a high mass resolution but also a sufficient temporal resolution.
@@
The timestep constraint of these simulations comes at a significant computational expense. Longer timesteps than the time-scale of cooling or adiabatic expansion can fail to estimate the internal energy of the gas. This is demonstrated in, e.g., Saitoh & Makino (2009) that longer time steps, which cannot properly resolve the evolution of expanding regions, lead to catastrophic results. Capturing the initial state of the energy injection requires a timescale of hundreds of years, which is three orders of magnitude shorter than the typical ISM in galaxy simulations. Proper timestep limiters are typically employed to preserve thermal energy injection by limiting the timestep with the signal velocity of a fluid tracer (Monaghan, 1997; Durier & Dalla Vecchia, 2012; Saitoh & Makino, 2009). Even colder particles should have a smaller timestep to prevent vastly different time steps between neighboring gas particles (e.g., Chaikin et al., 2023; Nobels et al., 2023). Hence, it might be desirable to speed up the injection scheme in a physically motivated manner.
@@
Machine learning (ML) may help to extend the range of physical scales in computationally expensive simulations (e.g., Kochkov et al., 2021). In cosmological simulations, ML has been used to predict the gravitational dynamics, allowing to avoid the computation of the physical states of $>10^{9}$ particles over the periods of the universe’s age (He et al., 2019; Jamieson et al., 2023, 2024). ML has also been incorporated into hydrodynamical simulations to reconstruct the physical states of fluid tracer particles undergoing complex astrophysical interactions in realistic systems (Hirashima et al., 2023a, b; Chan et al., 2024).
@@
An alternative is the data-driven approach, which studies spatial and temporal correlations in the distribution of physical data without prior knowledge. This approach has been applied for non-linear and multiscale physical simulations in astrophysics in recent work (He et al., 2019; Bernardini et al., 2022; Hirashima et al., 2023a, b; Jamieson et al., 2023, 2024; Chan et al., 2024; Legin et al., 2024). In this paper, we present the first surrogate modeling using ML to accelerate galaxy formation simulations by mitigating the bottlenecks caused by small timesteps.
@@ 2.1 Simulation code
This study executes three types of runs: high-resolution simulations of SN feedback in isolated molecular clouds, a fiducial run of an isolated dwarf galaxy, and a run of the same galaxy enhanced using the surrogate model for SN feedback. The simulation data of isolated molecular clouds are used to train the surrogate model. The isolated SN simulations and the fiducial run of the galaxy are carried out with our $N$-body/density-independent smoothed particle hydrodynamics (DISPH; Saitoh & Makino, 2013, 2016) code, ASURA-FDPS (Saitoh et al., 2008; Iwasawa et al., 2016; Hirashima et al., 2023a). Hydrodynamic interactions are computed by the DISPH method (Saitoh & Makino, 2013, 2016), which can properly handle the contact discontinuity. The gas properties are smoothed using the Wendland $C^{4}$ kernel (Dehnen & Aly, 2012). The kernel size is determined to maintain the number of neighbor particles at $125\pm 1$. We adopt the shared variable timesteps with the Courant-like hydrodynamical timestep (Courant et al., 1928) based on the signal velocity (Monaghan, 1997; Springel, 2005). Following the leapfrog (Velocity Verlet) scheme, we adopt a second-order symplectic integrator in its Kick-Drift-Kick (KDK) configuration. As for the simulations of SN feedback with a high-mass resolution, the timestep of a SPH particle $i$ is determined as
@@
$w_{ij}$ $={\bm{v}}_{ij}\cdot{\bm{r}}_{ij}/r_{ij},$ (2)
@@
Our new presenting framework ASURA-FDPS-ML consists of ASURA-FDPS and the surrogate modeling for SN feedback, which will be described in Sec. 2.6. As for dwarf galaxy simulations with surrogate modeling, timesteps are fixed to 2000 years by avoiding direct calculations of the initial state of SN feedback, which significantly requires smaller timesteps.
@@ 2.2 Initial Condition
The initial condition (IC) for the isolated galaxy simulations are originally described in Hu et al. (2016, 2017) (“cmp”). This IC has been used to investigate the properties of the dwarf galaxy and treatment of stellar feedback (e.g., Smith et al., 2021; Hislop et al., 2022; Steinwandel & Goldberg, 2023; Hu et al., 2023). The ICs were generated with the method proposed in Springel (2005). The dark matter halo follows a Hernquist profile with an NFW-equivalent (Navarro et al., 1997) concentration parameter $c_{\mathrm{NFW}}=10$ and has a virial radius of $R_{\mathrm{vir}}=44$ kpc and a viral mass of $M_{\mathrm{vir}}=2\times 10^{10}~{}\mathrm{M_{\odot}}$. The initial gas mass of the system is $4\times 10^{7}~{}\mathrm{M_{\odot}}$, and the stellar background potential has a mass of $2\times 10^{7}~{}\mathrm{M_{\odot}}$. The initial disk consists of 4 million dark matter particles, 10 million gas particles, and 5 million star particles, setting a dark matter particle mass resolution of $m_{\mathrm{DM}}=6.8\times 10^{3}~{}\mathrm{M_{\odot}}$ and a baryonic particle mass resolution of $m_{\mathrm{baryon}}=4~{}\mathrm{M_{\odot}}$. The gravitational softening parameter of the simulations is set to 62 pc for dark matter and 0.5 pc for gas and stars. Unlike some previous studies, the IC used in this paper has evolved in advance for 500 Myr with the code P-Gadget3 in the MFM version of Steinwandel et al. (2020) with 32 neighbors for the density computation. Fluid fluxes up to that point have been computed with the Harten-Lax-van-Leer contact (HLLC) Riemann solver. After this initial 500 Myr period, we completely switch to the ASURA-FDPS (or ASURA-FDPS-ML) framework based on the DISPH method outlined in Sec. 2.1.
@@ 2.3 Star formation model
Gas particles stochastically form stars if they satisfy several physical conditions such as high number density of hydrogen atoms ($n_{\mathrm{H}}\geq 100~{}\mathrm{cm^{-3}}$), low temperature ($T\leq 100$ K), and convergence of the flows ($\nabla\cdot v<0$), following Hirai et al. (2021). To avoid star formation in hot regions, which have a large Jeans mass, the threshold temperature 100 K is chosen following Hu et al. (2016). The masses are sampled by the initial mass function (IMF), following Chabrier (2003) with an upper limit of 40 M and a fixed star formation efficiency of 2 percent per free-fall time. When a gas particle meets the physical conditions, it is converted into a star with a probability ($p_{*}$) in a given timestep ($\Delta t$):
@@
where $m_{\mathrm{gas}}$, $\langle m_{*}\rangle$, $t_{\mathrm{dyn}}$, and $c_{*}$ are the mass of one gas particle, the average value of stellar mass in the assumed IMF, the dynamical time of the star-forming region, and a dimensionless star-formation efficiency.
@@ 2.4 Stellar feedback
We used the yields from Nomoto et al. (2013), which covers core-collapse SNe from 13 M to 40 M, where the nucleosynthesis yields is given as functions of the stellar mass, metallicity, and explosion energy of 1051 erg. These are compiled in the Chemical Evolution Library (CELib, Saitoh, 2017).
@@ 2.5 Required timestep for SN feedback
where $\rho$ is the density of a SPH particle. Let $N_{\mathrm{NB}}$ be the number of the neighboring particles that receive the energy from a SN. The internal energy for a SPH particle due to the SN of a single star, $U_{\mathrm{SN}}$, is defined as
@@
$\simeq\frac{5.0\times 10^{17}}{N_{\mathrm{NB}}}~{}\left(\frac{1~{}M_{\odot}}{m_{\mathrm{SPH}}}\right)~{}[\mathrm{erg~{}g^{-1}}],$ (5)
@@
$\simeq\frac{7.5\times 10^{3}}{N_{\mathrm{NB}}^{1/2}}~{}\left(\frac{1~{}M_{\odot}}{m_{\mathrm{SPH}}}\right)^{1/2}~{}[\mathrm{km~{}s^{-1}}].$ (6)
@@
$t_{\mathrm{SN}}\simeq 1.3\times 10^{3}\left(\frac{m_{\mathrm{SPH}}}{1~{}M_{\odot}}\right)^{5/6}\left(\frac{100~{}[\mathrm{cm^{-3}}]}{n_{\mathrm{H}}}\right)^{1/3}~{}[\mathrm{yr}].$ (7)
To avoid overcooling of SN feedback in star-forming regions, we replace the direct computation of SN feedback in denser regions than a hydrogen number density of $n_{\mathrm{H}}=1~{}\mathrm{cm^{-3}}$ with our surrogate model including ML. Using Equation (7), the timescale for SN feedback is calculated as $t_{\mathrm{SN}}\simeq 2\times 10^{4}$ years at a mean density of $n_{\mathrm{H}}=1~{}\mathrm{cm^{-3}}$ when the mass resolution $m_{\mathrm{SPH}}=4~{}M_{\odot}$ is applied. Given the typical Courant time-step in the region is the one-tenth of the timescale of SNe, we set a fixed timestep $\Delta t=0.1\,t_{\mathrm{SN}}=2\times 10^{3}$ years, sufficiently small to resolve SNe in the ISM with a density lower than $1~{}\mathrm{cm^{-3}}$.

### v2

@@ Abstract
We introduce new high-resolution galaxy simulations accelerated by a surrogate model that reduces the computation cost by approximately 75 percent. Massive stars with a Zero Age Main Sequence mass of more than about 10 $\mathrm{M_{\odot}}$ explode as core-collapse supernovae (CCSNe), which play a critical role in galaxy formation. The energy released by CCSNe is essential for regulating star formation and driving feedback processes in the interstellar medium (ISM). However, the short integration timesteps required for SNe feedback have presented significant bottlenecks in astrophysical simulations across various scales. Overcoming this challenge is crucial for enabling star-by-star galaxy simulations, which aim to capture the dynamics of individual stars and the inhomogeneous shell’s expansion within the turbulent ISM. To address this, our new framework combines direct numerical simulations and surrogate modeling, including machine learning and Gibbs sampling. The star formation history and the time evolution of outflow rates in the galaxy match those obtained from resolved direct numerical simulations. Our new approach achieves high-resolution fidelity while reducing computational costs, effectively bridging the physical scale gap and enabling multi-scale simulations.
@@
thanks: JSPS Research Fellow
@@ 1 Introduction
Supernovae (SNe) are powerful events that release an immense amount of energy into the ambient ISM ($\sim 10^{51}$ erg). This energy release results in immediate heating and delayed momentum input into the ambient ISM, driving turbulence and aiding star formation regulation in the turbulent ISM (e.g., Somerville & Davé, 2015; Naab & Ostriker, 2017, for reviews). While SNe strongly dominate the energy budget for galaxy formation and evolution, the evolution of galaxies also involves complex interactions among gravity, hydrodynamics, radiation, star formation, and chemical reactions. Given the complicated interplay among these processes, numerical methods have been commonly employed to study galaxy formation and evolution.
Recently, numerical simulations have reached the resolutions of $\sim 1~{}\mathrm{M_{\odot}}$ (e.g., Andersson et al., 2023, 2024; Deng et al., 2024; Emerick et al., 2018, 2019; Fujii et al., 2021, 2021, 2022b, 2022a, 2024; Gutcke et al., 2022a, b; Hirai et al., 2021, 2025; Hu et al., 2016, 2017; Hu, 2019; Lahén et al., 2020, 2023; Smith et al., 2021; Steinwandel et al., 2020; Steinwandel & Goldberg, 2025; Steinwandel et al., 2023; Lahén et al., 2025; Steinwandel et al., 2024a, b), which allow to resolve several important phases of SN-remnant evolution including the energy-conserving Sedov-Taylor phase, the pressure-driven snowplow phase, and the momentum-conserving snowplow phase (e.g., Sedov, 1959; Taylor, 1950; Blondin et al., 1998; von Neumann, 1942) from first principles.
@@
The timestep constraint of these simulations comes at a significant computational expense. Longer timesteps than the time-scale of cooling or adiabatic expansion tend to fail to estimate the internal energy of the gas and could lead to catastrophic results (e.g., Saitoh & Makino, 2009). Capturing the Sedov phase, the initial evolution phase of explosions just after the energy injection, requires a timescale of hundreds of years, three orders of magnitude shorter than that of the typical ISM in galaxy simulations. Proper timestep limiters are typically employed to preserve thermal energy injection by limiting the timestep with the signal velocity of a fluid tracer (Monaghan, 1997; Durier & Dalla Vecchia, 2012; Saitoh & Makino, 2009). Even colder smoothed particle hydrodynamics (SPH) particles should have a smaller timestep to prevent vastly different time steps between neighboring SPH particles (e.g., Chaikin et al., 2023; Nobels et al., 2024). Hence, it might be desirable to speed up the injection scheme in a physically motivated manner.
@@
Machine learning (ML) may help to extend the range of physical scales in computationally expensive simulations (e.g., Kochkov et al., 2021). In cosmological simulations, ML has been used to predict the gravitational dynamics, allowing to avoid the computation of the physical states of $>10^{9}$ particles over the periods of the universe’s age (He et al., 2019; Jamieson et al., 2023, 2025). ML has also been incorporated into hydrodynamical simulations to reconstruct the physical states of fluid tracer particles undergoing complex astrophysical interactions in realistic systems (Hirashima et al., 2023a, b; Chan et al., 2024).
@@
An alternative is the data-driven approach, which studies spatial and temporal correlations in the distribution of physical data without prior knowledge. This approach has been applied for non-linear and multiscale physical simulations in astrophysics in recent work (He et al., 2019; Bernardini et al., 2022; Hirashima et al., 2023a, b; Jamieson et al., 2023, 2025; Chan et al., 2024; Legin et al., 2024). In this paper, instead of PINNs, we present the first data-driven surrogate modeling using ML to accelerate galaxy formation simulations by mitigating the bottlenecks caused by small timesteps.
@@ 2.1 Simulation code
This study executes three types of runs: high-resolution simulations of SN feedback in isolated molecular clouds, a fiducial run of an isolated dwarf galaxy, and a run of the same galaxy enhanced using the surrogate model for SN feedback. The simulation data of isolated molecular clouds are used to train the surrogate model. The isolated SN simulations and the fiducial run of the galaxy are carried out with our $N$-body/density-independent smoothed particle hydrodynamics (DISPH; Saitoh & Makino, 2013, 2016) code, ASURA-FDPS (Saitoh et al., 2008; Iwasawa et al., 2016; Hirashima et al., 2023a). Hydrodynamic interactions are computed by the DISPH method (Saitoh & Makino, 2013, 2016), which can properly handle the contact discontinuity. The gas properties are smoothed using the Wendland $C^{4}$ kernel (Dehnen & Aly, 2012). The kernel size is determined to maintain the number of neighbor particles at $125\pm 1$. We adopt the shared variable timesteps with the Courant-like hydrodynamical timestep (Courant et al., 1928) based on the signal velocity (Monaghan, 1997; Springel, 2005). Following the leapfrog (Velocity Verlet) scheme, we adopt a second-order symplectic integrator in its Kick-Drift-Kick (KDK) configuration. In the simulations of SN feedback with a high-mass resolution, the timestep of a SPH particle $i$ is determined as
@@
$w_{ij}=\begin{cases}{\bm{v}}_{ij}\cdot{\bm{r}}_{ij}/r_{ij}&({\bm{v}}_{ij}\cdot{\bm{r}}_{ij}<0),\\ 0,&({\bm{v}}_{ij}\cdot{\bm{r}}_{ij}\geq 0).\end{cases}$ (2)
@@
Our new presenting framework ASURA-FDPS-ML consists of ASURA-FDPS and the surrogate modeling for SN feedback, which will be described in Sec. 2.6. In our dwarf galaxy simulations with surrogate modeling, we fix timesteps to 2000 years by avoiding direct calculations of the initial state of SN feedback, which significantly requires smaller timesteps.
@@ 2.2 Initial Condition
The initial condition (IC) for the isolated galaxy simulations are originally described in Hu et al. (2016, 2017) (“cmp”). This IC has been used to investigate the properties of the dwarf galaxy and treatment of stellar feedback (e.g., Smith et al., 2021; Hislop et al., 2022; Steinwandel & Goldberg, 2025; Hu et al., 2023). The ICs were generated with the method proposed in Springel (2005). The dark matter halo follows a Hernquist profile with an NFW-equivalent (Navarro et al., 1997) concentration parameter $c_{\mathrm{NFW}}=10$ and has a virial radius of $R_{\mathrm{vir}}=44$ kpc and a viral mass of $M_{\mathrm{vir}}=2\times 10^{10}~{}\mathrm{M_{\odot}}$. The initial gas mass of the system is $4\times 10^{7}~{}\mathrm{M_{\odot}}$, and the stellar background potential has a mass of $2\times 10^{7}~{}\mathrm{M_{\odot}}$. The initial disk consists of 4 million dark matter particles, 10 million gas particles, and 5 million star particles, setting a dark matter particle mass resolution of $m_{\mathrm{DM}}=6.8\times 10^{3}~{}\mathrm{M_{\odot}}$ and a baryonic particle mass resolution of $m_{\mathrm{baryon}}=4~{}\mathrm{M_{\odot}}$. The gravitational softening parameter of the simulations is set to 62 pc for dark matter and 0.5 pc for gas and stars. Unlike some previous studies, the IC used in this paper has evolved in advance for 500 Myr with the code P-Gadget3 in the MFM version of Steinwandel et al. (2020) with 32 neighbors for the density computation. Fluid fluxes up to that point have been computed with the Harten-Lax-van-Leer contact (HLLC) Riemann solver. After the initial 500 Myr period, we completely switch to the ASURA-FDPS (or ASURA-FDPS-ML) framework based on the DISPH method outlined in Sec. 2.1. We note that only a small amount of gas is converted into stars during the pre-evolution phase, and, for simplicity, we ignore them in the subsequent evolution. See Appendix B for further discussion on the switching effects. We adopt an initial metallicity of 0.1 $\mathrm{Z_{\odot}}$ following the solar abundance pattern and only use the stellar background particles, excluding massive stars, when switching the framework.
@@ 2.3 Star formation model
Gas particles stochastically form stars if they satisfy several physical conditions such as high number density of hydrogen atoms ($n_{\mathrm{H}}\geq 100~{}\mathrm{cm^{-3}}$), low temperature ($T\leq 100$ K), and convergence of the flows ($\nabla\cdot v<0$), following Hirai et al. (2021). To avoid star formation in hot regions, which have a large Jeans mass, the threshold temperature 100 K is chosen following Hu et al. (2016). The masses are sampled by the initial mass function (IMF), following Chabrier (2003) with an upper limit of 40 M. The lower limit for the IMF is 0.1 M. When a gas particle meets the physical conditions, it is converted into a star with a probability ($p_{*}$) in a given timestep ($\Delta t$):
@@
where $m_{\mathrm{gas}}$, $\langle m_{*}\rangle$, $t_{\mathrm{dyn}}=1/\sqrt{4\pi G\rho}$, and $c_{*}$ are the mass of one gas particle, the average value of stellar mass in the assumed IMF, the local dynamical time of the star-forming region, and a fixed dimensionless star-formation efficiency as 0.02 per dynamical time (Krumholz et al., 2019).
@@ 2.4 Stellar feedback
We used the yields from Nomoto et al. (2013), which covers core-collapse SNe from 13 M to 40 M, where the nucleosynthesis yields as masses of individual elements are given as functions of the stellar mass, metallicity, and explosion energy of 1051 erg. The lifetimes are determined depending on the metallicity using the table in Portinari et al. (1998). These are compiled in the Chemical Evolution Library (CELib, Saitoh, 2017).
@@ 2.5 Required timestep for SN feedback
where $\rho$ is the density of a SPH particle. Let $N_{\mathrm{NB}}$ be the number of neighboring particles that receive energy from a SN. The internal energy for a SPH particle due to the SN of a single star, $U_{\mathrm{SN}}$, is defined as
@@
$\simeq\frac{5.0\times 10^{17}}{N_{\mathrm{NB}}}~{}\left(\frac{1~{}\mathrm{M_{\odot}}}{m_{\mathrm{SPH}}}\right)~{}[\mathrm{erg~{}g^{-1}}],$ (5)
@@
$\simeq\frac{7.5\times 10^{3}}{N_{\mathrm{NB}}^{1/2}}~{}\left(\frac{1~{}\mathrm{M_{\odot}}}{m_{\mathrm{SPH}}}\right)^{1/2}~{}[\mathrm{km~{}s^{-1}}].$ (6)
@@
$t_{\mathrm{SN}}\simeq 1.3\times 10^{3}\left(\frac{m_{\mathrm{SPH}}}{1~{}\mathrm{M_{\odot}}}\right)^{5/6}\left(\frac{100~{}[\mathrm{cm^{-3}}]}{n_{\mathrm{H}}}\right)^{1/3}~{}[\mathrm{yr}].$ (7)
To avoid overcooling of SN feedback in star-forming regions, we replace the direct computation of SN feedback in denser regions than a hydrogen number density of $n_{\mathrm{H}}=1~{}\mathrm{cm^{-3}}$ with our surrogate model including ML. Using Equation (7), the timescale for SN feedback is calculated as $t_{\mathrm{SN}}\simeq 2\times 10^{4}$ years at a mean density of $n_{\mathrm{H}}=1~{}\mathrm{cm^{-3}}$ when the mass resolution $m_{\mathrm{SPH}}=4~{}M_{\odot}$ is applied. Given that the typical Courant time-step in the region is one-tenth of the timescale of SNe, we set a fixed timestep $\Delta t=0.1\,t_{\mathrm{SN}}=2\times 10^{3}$ years, sufficiently small to resolve SNe in the ISM with a density lower than $1~{}\mathrm{cm^{-3}}$.
We also note that, in simulations with a fixed timestep in our new framework, the Courant condition of equation (1) is satisfied for almost all particles. On average, only 0.08 particles (0.32 $\mathrm{M_{\odot}}$) per snapshot require timesteps shorter than the fixed timestep of 2000. Most of these particles require approximately 1900 years, with the shortest requiring about 1400 years. This difference is less than 30 percent smaller than the fixed timestep, and since these shorter timesteps are conservatively estimated with the relative velocity as $3w_{ij}$ using equation (1), such minor differences are not significant. Therefore, we conclude that while a tiny number of particles technically violate the Courant condition, their impact on the simulation results is negligible due to both their small number and the minor degree of violation.